OpenStudy (anonymous):

Find the missing value. I = $95.00 P = ? r = 5% t = 4 years $0.21 $475.00 $237.50 $4.75

OpenStudy (anonymous):

I = P * R * T (The formula is used to calculate simple interest) I mostly need help checking my work. :P

OpenStudy (anonymous):

I = P * R * T 95 = P * 0.05 * 4 (95/ 0.05 = 1900) 1900 = P *4 (1900/4 = 475) 475 = P Therefore, the answer is B... correct?

OpenStudy (freckles):

\[I=P R T \implies P=\frac{I}{RT} \\ P=\frac{95}{.05 \cdot 4}=\frac{95}{.2}\] your answer is right

OpenStudy (freckles):

great job and i think it is totally cool you showed your work go you! :)

OpenStudy (anonymous):

Aye, thanks. I'm never too confident in myself. :P Have a great day, dude.

OpenStudy (freckles):

and your work does look awesome by the way every step is done really well

OpenStudy (anonymous):

5000 = P x 0.025 x 1 (5000/ 0.025 = 200000) 200000 = P x 1 200000 = P ?

OpenStudy (anonymous):

Just a follow up question. ;P